Two dice are thrown. Find the probability that the number appeared have a sum 8 if it is known
that second dice always exhibits 4.

Grade:Upto college level

3 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Second die always exhibits 4 so (1,4)(2,4)(3,4)(4,4)(5,4)(6,4)
as second die is always 4 and to sum 8 (4,4) is only possibility
P(B/A)=1/6
Thanks and Regards
Shaik Aasif
askIITians faculty
ANGEL MARY THOMAS
6 years ago
P(S)=36
Event A :-Numbers appeared have a sum 8
              {(2,6),(6,2),(3,5),(5,3),(4,4)}
Event B:-second die always exhibit 4
              {(1,4),(2,4),(3,4),(4,4),(5,4),(6,4)}
A INTERSECTION B={(4,4)}
P(B/A)= [P(B INTERSECTION A)]/P(B)
             =[1/36 ]  / [6/36] 
               =1/6
Vikash
5 years ago
By conditional probability:
We konw that P(A/B)={P(A intersection b)}/P(B).
 
Here P(a)=no of event in which sum appears 8
P(B)= second die exhibit 4
 
 
P(A)=(2,6)(3,5)(4,4)5,3)(6,2)
 
P(B)=(1,4)(2,4)(3,4)(4,4)(5,4)(6,4)
 
 
P(A intersection B)=(4,4)
•°• P(B/A)=(1/36)/(6/36)=1/6